Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Race Probablility

Expert replies
by durgesh79 » Sat Jun 28, 2008 9:26 am
If X and Y are two of 5 participants in a race, then what is the probability that X will always beat Y.
Join the discussion
Source: — Problem Solving |

by spanlength » Sat Jun 28, 2008 10:08 am
So x can get a position from 1 to 4 in race.
Hece the probablitity =x is 1st in race * y can get 2nd,3rd , 4th or 5th +x is 2nd in race y can get 3rd or 4th, 5th + 3rd in race * y can be or 4th or 5th + 4th in race * ycan be 5th in race = 1/5*4/5 +1/5* 3/5 +1/5* 2/5+1/5*1/5 =2/5.
Last edited by spanlength on Sat Jun 28, 2008 10:21 am, edited 1 time in total.
Join the discussion

by antyagi » Sat Jun 28, 2008 10:16 am
But then the probablity of Y also needs to be linked.
IF X is 1st Y can be 2,3,4,5
IF X is 2 Y can be 3,4,5

Hence it would be : 1/5 * 4/5 + 1/5*3/5 + 1/5 *2/5 + ....
Thanks,
Ankur Tyagi
Join the discussion

by durgesh79 » Sat Jun 28, 2008 11:01 am
2/5 is not the answer
Join the discussion

by gabriel » Sat Jun 28, 2008 11:27 am
It would be true half of the time. So probability is 1/2
Join the discussion

by Ian Stewart » Sat Jun 28, 2008 11:35 am
I assume the order of finish is random? Then it's 50% that x finishes ahead of y, and 50% that y finishes ahead of x.
Join the discussion

by durgesh79 » Sat Jun 28, 2008 12:36 pm
Right. The answer is 1/2

If X is on 1, then y can take any 4 of the 4 places left -> 1/5 * 4/4
If X is on 2, then y can take any 3 of the 4 places left -> 1/5 * 3/4
If X is on 3, then y can take any 2 of the 4 places left -> 1/5 * 2/4
If X is on 4, then y can take only 1 of the 4 places ledt -> 1/5 * 1/4

Probability = sum of all these = 1/5 * ( 1+3/4+1/2+1/4)
= 1/2

Buti knew that there could be a better way to do it.
Join the discussion

Why

by evansbd » Thu Jul 17, 2008 8:45 am
Ian

Intuitively, can you explain how you know that X always beats Y and vice versa are 50%? If its a two person race, then it would be obvious, but I dont see how that could jump out for a race such as in the question.

Is there a way to do this problem using comibinations?

i.e. There are 5! possibilities for a finishing order, only a percentage of these will result in X beating Y...

Also, if there were more participants, and we were still concerned with X always beating Y, would the answer still be 1/2?
Join the discussion

Why

by evansbd » Fri Jul 18, 2008 6:09 am
repost
Join the discussion

by pepeprepa » Fri Jul 18, 2008 7:39 am
I give you two explanations.

If you do not see it, take an example (with 3 runners X,Y and A)
You want to know who wins most of the time between X and Y
Write all the possible results:
XAY XYA AXY
AYX YAX YXA
You can clearly see that the probability "X is before Y" is 1/2.
You can choose the "n" number of runners you want, it will be the same.

Another explanation:
« A »= A is before B
« B »= B is before A
P(A)+P(B)=1 and so P(A)=1/2 because
*)A cannot have the same rank as B in the race so the sum of the 2 probabilities is 1
*)There is equiprobability of both, noone is doping or cheat more than the other one, so they have the same probability to occur each other. Given the result is 1 we have 2P(A)=1
Join the discussion